OpenStudy (anonymous):

How many grams of acetic anhydride are needed to react with 5.46 g of salicylic acid?

OpenStudy (anonymous):

How many grams of aspirin will result?

OpenStudy (anonymous):

How many grams of acetic acid are formed as a by-product?

OpenStudy (koikkara):

Hint: Q1 ) molecular weight of salicylic acid\(=\)138.12 g/mol molecular weight of acetic anhydride \(=\) 102.09 g/mol. 5.40 g salicylic acid \(=\) 0.0391 moles. which means reaction is 1:1 ratio.. \(=\)---- g Q2 ) The molecular weight of aspirin \(=\) 180.157 g/mol. 0.0391 moles of this would be ________ g ? (approx 4.00) Q3) Molecular weight of acetic acid \(=\) 60.05 gm/mol, then______ g is formed. @mtuck

OpenStudy (photon336):

@mtuck start with the balanced chemical reaction first.

OpenStudy (anonymous):

i dont know how to balance

OpenStudy (anonymous):

@koikkara pleae help

OpenStudy (anonymous):

@Photon336 help please

OpenStudy (anonymous):

Aspirin is prepared by reaction of salicylic acid (C7H6O3) with acetic anhydride (C4H6O3) according to the following equation: C7H6O3Salicylicacid+C4H6O3Aceticanhydride→ C9H8O4Aspirin+CH3COOHAceticacid

OpenStudy (photon336):

so @mtuck this is our reaction right? C7H6O3 + C4H6O3 --> C9H8O4 + C2H4O2 so we know that one mole of salicylic acid reacts with 1 mol of acetic anhydride. our next step is to look up the molar masses of each of these compounds and then figure out how many moles we need. I'll show you how to do that.

OpenStudy (anonymous):

i finished it already :) thanks.

OpenStudy (anonymous):

it was due at 9:20

OpenStudy (photon336):

Could you show me what answer you got?

OpenStudy (anonymous):

#1 = 4.04 g

OpenStudy (photon336):

Ah okay, no problem one thing you should remember @mtuck is that to do stiochiometry you need to start with the balanced chemical reaction first. then you can convert grams to moles and son on

OpenStudy (photon336):

102.9 g/mol acetic anhydride 138.21g/mol salic acid \[(5.46grams) C_{7}H_{6}O_{3}*\frac{ mol }{ 138.21grams } = 0.0395 mol,C_{7}H_{6}O_{3}\] \[0.0395.mol C_{7}H_{6}O_{3}*\frac{ C_{4}H_{6}O_{3} }{ C_{7}H_{6}O_{3} } = 0.0395 mol, C_{4}H_{6}O_{3}*(\frac{ 102.9grams }{ mol} =4.06grams C_{4}H_{6}O_{3} \] nicely done